### v5

# On Minimal Depth in Neural Networks

Abstract
Understanding the relationship between the depth of a neural network and its representational capacity is a central problem in deep learning theory. In this work, we develop a geometric framework to analyze the expressivity of ReLU networks with the notion of depth complexity for convex polytopes. The depth of a polytope recursively quantifies the number of alternating convex hull and Minkowski sum operations required to construct it. This geometric perspective serves as a rigorous tool for deriving depth lower bounds and understanding the structural limits of deep neural architectures.
We establish lower and upper bounds on the depth of polytopes, as well as tight bounds for classical families. These results yield two main consequences. First, we provide a purely geometric proof of the expressivity bound by Arora et al. (2018), confirming that $\lceil\log_{2}(n+1)\rceil$ hidden layers suffice to represent any continuous piecewise linear (CPWL) function. Second, we prove that, unlike general ReLU networks, convex polytopes do not admit a universal depth bound. Specifically, the depth of cyclic polytopes in dimensions $n\geq 4$ grows unboundedly with the number of vertices. This result implies that Input Convex Neural Networks (ICNNs) cannot represent all convex CPWL functions with a fixed depth, revealing a sharp separation in expressivity between ICNNs and standard ReLU networks.
1 Introduction
A ReLU neural network of depth $m$, or with $m$ hidden layers, is a function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ expressed as
$f=T^{(m+1)}\circ\sigma_{m}\circ T^{(m)}\circ\sigma_{m-1}\circ\dots\circ T^{(2)}\circ\sigma_{1}\circ T^{(1)},$
where $T^{(l)}:\mathbb{R}^{n_{l-1}}\rightarrow\mathbb{R}^{n_{l}}$ are $m+1$ affine transformations, $n_{0}=n$ and $n_{m+1}=1$. The activation function $\sigma_{l}$ corresponds to the vectorized ReLU function $\sigma_{l}(x_{1},\ldots,x_{n_{l}})=(\max\{x_{1},0\},\max\{x_{2},0\},\ldots,\max\{x_{n_{l}},0\})$.
ReLU neural networks are continuous piecewise linear (CPWL) functions [2, 11]. An important problem in learning theory is to characterize which CPWL functions can be represented by a ReLU network of a given depth. Such expressivity results are relevant for quantifying depth-dependent convergence rates for various function classes [9, 10, 11, 25] and enhancing our theoretical understanding of deep architectures [2, 22, 12, 19].
We denote the set of all CPWL functions from $\mathbb{R}^{n}\rightarrow\mathbb{R}$ as $\mathsf{CPWL}_{n}$, and the subset of these functions that can be represented by ReLU networks of depth $m$ as $\mathsf{ReLU}_{n}(m)$.
It has been shown that $\lceil\log_{2}(n+1)\rceil$ hidden layers are sufficient to represent any CPWL function.
Theorem 1 (Arora, Basu, Mianjy and Mukherjee [2], Theorem 2.1).
$\mathsf{CPWL}_{n}=\mathsf{ReLU}_{n}(\lceil\log_{2}(n+1)\rceil).$
The minimal depth required to represent any CWPL function remains an open question. Hertrich, Basu, Di Summa, and Skutella [17] proposed an approach that reduces the problem to analyzing a single function.
Proposition 1 (Hertrich, Basu, Di Summa and Skutella [17], Proposition 1.6).
$\mathsf{CPWL}_{n}=\mathsf{ReLU}_{n}(m)\Longleftrightarrow\max\{x_{1},x_{2},\ldots,x_{n},0\}\in\mathsf{ReLU}_{n}(m).$
Following Proposition 1, the goal is to find the minimum $m$ such that the function $\max\{x_{1},\ldots,x_{n},0\}\in\mathsf{ReLU}_{n}(m)$.
Hertrich, Basu, Di Summa, and Skutella [17] also conjectured that this depth is exactly $\lceil\log_{2}(n+1)\rceil$, which had been proven for dimensions $n\leq 3$ [11, 21]. Much of the progress toward resolving this problem has come from formulating Proposition 1 in the language of convex polytopes, i.e. the convex hull of finitely many points [15, 28]. Haase, Hertrich, and Loho [16] proved the conjecture for ReLU networks with integer weights. This result was later extended by Averkov, Hojny, and Merkert [3] to networks with decimal fractional weights. Furthermore, Grillo, Hertrich, and Loho [14] proved the conjecture for $n=4$ in networks compatible with specific polyhedral complexes. More recently, the conjecture was disproved by Bakaev, Brunck, Hertrich, Stade, and Yehudayoff [5] via polyhedral subdivisions of the simplex; they showed that $\lceil\log_{3}(n-1)\rceil+1$ hidden layers are sufficient to represent any CPWL function.
In this work, we introduce the depth complexity of polytopes, a geometric analogue of neural network depth, to analyze the expressivity of ReLU networks. Although this concept has been used implicitly in prior work [17, 16, 5], we provide a systematic study of it.
We denote the depth complexity of a polytope $P$ as $d(P)$ and define it recursively as follows. If $P$ consists of a single point, we define $d(P)=0$. If $P$ is not a single point, we define $d(P)=m$, where $m$ is the smallest positive integer such that
$P=\sum_{i=1}^{q}\mathrm{conv}(P_{i1},P_{i2}),\quad\text{with }d(P_{ij})<m\quad\forall i,j.$ (1)
Here, $\mathrm{conv}(P_{1},P_{2})=\mathrm{conv}(P_{1}\cup P_{2})$ and $P_{1}+P_{2}=\{a_{1}+a_{2}:a_{i}\in P_{i}\}$ are the convex hull and Minkowski sum operations respectively.
At its core, the depth complexity of a polytope quantifies the number of alternating steps of convex hulls and Minkowski sums required to construct it. This serves as a measure of how “complex” it is to represent a polytope using these two operations. Under this measure, the simplest polytope is a point and depth-1 polytopes correspond to zonotopes, which are Minkowski sums of segments.
Before connecting the depth of polytopes and ReLU networks, we introduce some terminology. A linear max function is a function of the form $f(x)=\max\{a_{1}\cdot x,\ldots,a_{p}\cdot x\}$ with $a_{i}\in\mathbb{R}^{n}$ for all $i$. Notably, linear max functions are convex and positively homogeneous, meaning they satisfy $f(\lambda x)=\lambda f(x)$ for $\lambda\geq 0$.
Let $\mathsf{LM}_{n}$ denote the collection of linear max functions from $\mathbb{R}^{n}\rightarrow\mathbb{R}$, and let $\mathcal{P}_{n}$ represent the set of convex polytopes in $\mathbb{R}^{n}$. These two sets constitute semirings with the max and sum operations, and convex hull and Minkowski sum operations, respectively.
For the function $f(x)=\max\{a_{1}\cdot x,\ldots,a_{p}\cdot x\}$, its Newton polytope is defined as
$\mathcal{N}f:=\mathrm{conv}(a_{1},\ldots,a_{p}).$
Similarly, for the polytope $P=\mathrm{conv}(a_{1},\ldots,a_{p})$, the associated support function is
$\mathcal{F}P(x):=\max\{a_{1}\cdot x,\ldots,a_{p}\cdot x\}.$
The mappings $\mathcal{N}$ and $\mathcal{F}$ are isomorphisms between the semirings ($\mathsf{LM}_{n}$, max, ++) and ($\mathcal{P}_{n}$, conv, ++) [20, 27].
The following result establishes a depth connection between ReLU networks and polytopes.
Theorem 2 (Hertrich, Basu, Di Summa and Skutella [17], Theorem 5.2).
Let $f\in\mathsf{CPWL}_{n}$ be a positively homogeneous function. Then, $f\in\mathsf{ReLU}_{n}(m)$ if and only if there exist $f_{i}\in\mathsf{LM}_{n},i=1,2,$ such that $f=f_{1}-f_{2}$ and $d(\mathcal{N}f_{i})\leq m$.
Combining Proposition 1 with Theorem 2 gives an equivalent geometric statement of Proposition 1.
Theorem 3.
$\mathsf{CPWL}_{n}=\mathsf{ReLU}_{n}(m)$
if and only if there exist polytopes $P_{i},i=1,2,$ such that
$P_{1}=\mathcal{N}(\max\{x_{1},x_{2},\ldots,x_{n},0\})+P_{2}\quad\text{and}\quad d(P_{i})\leq m.$
To fully leverage Theorem 3, it is relevant to establish foundational results on the depth complexity of polytopes. In addition, two natural questions arise: What is the minimum depth required to represent any polytope? And, what is the depth complexity of simplices like $\mathcal{N}(\max\{x_{1},x_{2},\ldots,x_{n},0\})$?
Depth Complexity Results
In Section 2, we derive depth complexity upper bounds based on Minkowski sums, affine transformations, convex hull representations, and combinatorial data such as the number of vertices, edges, and 2-faces. Combined with Theorem 2, these bounds yield depth estimates for ReLU network representations of linear max functions.
An affine transformation is a mapping $\varphi(x)=Ax+b$, where $A$ is a linear map and $b$ is a translation vector. A convex hull representation of a polytope $P$ in terms of other polytopes $P_{1},\ldots,P_{k}$ is a decomposition of the form $P=\mathrm{conv}(P_{1},P_{2},\ldots,P_{k})$.
If the polytopes $P_{i}$ are the vertices, edges or 2-faces of $P$, we can compute depth estimates based on their counts. For a polytope with $k$ vertices, we obtain the upper bound $\lceil\log_{2}k\rceil$, which is tight for certain families of polytopes.
In Section 3, we derive lower bounds on depth complexity using its graph and face structure. Let $G(P)$ denote the 1-skeleton of $P$, i.e. the graph whose vertices are those of $P$ and in which two vertices are adjacent when they span an edge of $P$.
We show that if $G(P)$ contains a complete subgraph on $k$ vertices, then $d(P)\geq\lceil\log_{2}k\rceil$. This follows from the fact that complete subgraphs are pervasive in expressions like (1): if $G(P)$ contains a complete subgraph, then at least a summand in the Minkowski sum will have one, which then propagates through the polytopes in the convex hull.
Moreover, for any face $F$ of $P$, we have $d(F)\leq d(P)$. Hence, $d(P)$ is at least the maximum depth complexity among its faces.
In Section 4, we determine the depth complexity of several families of polytopes, including pyramids, bipyramids, prisms, simplices, and cyclic polytopes. Using depth estimates based on vertex counts and complete subgraphs, it follows directly that polytopes with $k$ vertices whose graphs are complete, such as simplices and cyclic polytopes, have depth complexity $\lceil\log_{2}k\rceil$.
This result answers both motivating questions posed earlier. For simplices, the main application follows from Theorem 3, where we recover the same bound as in Theorem 1, thus providing an alternative geometric proof. In contrast, for dimension $n\geq 4$, cyclic polytopes form a family whose depth grows unbounded as the number of vertices increases. This shows that there is no universal upper depth bound to represent polytopes–unlike the case for ReLU networks in Theorem 1. Notably, the proofs of these results rely only on elementary polyhedral theory.
We conclude Section 4 by showing that cyclic polytopes not only exhibit unbounded depth, but in dimensions $n\geq 5$, their Minkowski sum with zonotopes yield, for any given depth $m$, families of polytopes with an increasing number of vertices, all of depth $m$.
Implications for Input Convex Neural Networks
Input Convex Neural Networks (ICNNs) [1] are ReLU networks constrained to use only monotone affine transformations and skip connections from the input layer. Here, a monotone affine transformation means an affine map $Ax+b$ where the matrix $A$ has nonnegative entries.
ICNNs are interesting because they can represent any convex CPWL function [4, 13] and have found many successful applications [18, 24, 6, 7].
Following [4], we can define under the ICNN model, a corresponding depth complexity $d_{0}(P)$ for a polytope $P$: $d_{0}(P)=0$ if $P$ is a single point, else $d_{0}(P)=m$ for $m$ the minimum value such that
$P=\sum_{i=1}^{p}\mathrm{conv}(P_{i1},P_{i2}),\quad\text{with }d_{0}(P_{i1})<m\text{ and }d_{0}(P_{i2})=0\quad\forall i.$
In [4] the authors showed that $d_{0}(\mathcal{N}(\max\{x_{1},\ldots,x_{n}\}))=n$, and that for $n=3$ there exists a family of polytopes with unbounded $d_{0}$-complexity. For ICNNs, $d_{0}(P)$ also constitutes the minimum ICNN depth required to represent $\mathcal{F}P$.
Since $d(P)\leq d_{0}(P)$, and cyclic polytopes with $k$ vertices satisfy $d(P)=\lceil\log_{2}k\rceil$ for $n\geq 4$, they also form an unbounded depth family for $d_{0}$. Thus, although ICNNs can represent every convex CPWL function, no fixed depth bound suffices in general. This contrasts with the universal depth bound for ReLU networks established in Theorem 1.
Acknowledgements. I am deeply grateful to Francisco Santos for Theorem 5, valuable discussions, and his hospitality at the University of Cantabria. I also thank Ansgar Freyer for discussions on the depth complexity of simplices.
2 Depth Upper Bounds
We begin by computing basic bounds for Minkowski sums, convex hulls and affine maps.
Proposition 2.
Let $P_{i},i=1,2,$ be polytopes with $d(P_{i})\leq m_{i}$ and $\varphi$ be an affine transformation. Then,
  1. (a)
    $d(P_{1}+P_{2})\leq\max\{m_{1},m_{2}\}$
  2. (b)
    $d(\mathrm{conv}(P_{1},P_{2}))\leq\max\{m_{1},m_{2}\}+1$
  3. (c)
    $d(\varphi(P_{1}))\leq d(P_{1})$
Proof.
These depth bounds follow immediately from the depth complexity definition. If $d(P_{i})\leq m_{i}$, it implies that we can express $P_{i}$ as
$P_{i}=\sum_{j=1}^{q_{i}}\mathrm{conv}(Q_{j,i},R_{j,i}),\quad d(Q_{j,i}),d(R_{j,i})<m_{i}\quad\forall i,j.$
Then,
$P_{1}+P_{2}=\sum_{j=1}^{q_{1}}\mathrm{conv}(Q_{j,1},R_{j,1})+\sum_{j=1}^{q_{2}}\mathrm{conv}(Q_{j,2},R_{j,2}).$
All $Q_{j,i}$ and $P_{j,i}$ have depth strictly less than $\max\{m_{1},m_{2}\}$, therefore $d(P_{1}+P_{2})\leq\max\{m_{1},m_{2}\}$.
Similarly, since $d(P_{i})\leq\max\{m_{1},m_{2}\}$, it follows by definition that
$d(\mathrm{conv}(P_{1},P_{2}))\leq\max\{m_{1},m_{2}\}+1.$
Let $\varphi(x)=Ax+b$ be an affine map, if $d(P_{1})=0$, it is trivial that $d(\varphi(P_{1}))=0$. For the purpose of induction, assume the statement is true up to m1m-1. If $d(P_{1})=m$, then
$\varphi(P_{1})=\varphi\Big(\sum_{j=1}^{q_{1}}\mathrm{conv}(Q_{j,1},R_{j,1})\Big)=\sum_{j=1}^{q_{1}}\mathrm{conv}(AQ_{j,1},AR_{j,1})+b,$
and by induction hypothesis $d(AQ_{j,1}),d(AR_{j,1})<m$, implying $d(\varphi(P_{1}))\leq d(P_{1})$. ∎
If an affine transformation $\varphi$ is invertible, note that $d(\varphi(P))=d(P)$. This also implies that $d(P_{1}+P_{2})=d(P_{1})$ for $d(P_{2})=0$, which is simply a translation.
When $P=\mathrm{conv}(P_{1},\ldots,P_{k})$, iterating Proposition 2(b) yields an upper bound for $d(P)$. There are several ways to iterate a binary operation on $k$ elements, corresponding to different choices of binary bracketing.
Any binary bracketing of $\mathrm{conv}(P_{1},\ldots,P_{k})$ can be encoded by a full binary tree whose leaves are labeled by $1,\ldots,k$: each internal node records one application of $\mathrm{conv}(\cdot,\cdot)$, and its two children correspond to the two subexpressions being combined. Conversely, every full binary tree with leaves labeled by $1,\ldots,k$ determines a binary bracketing of $\mathrm{conv}(P_{1},\ldots,P_{k})$. For a leaf $i$, let $\ell_{i}$ denote its depth in the tree, that is, its distance to the root.
We identify the smallest bound obtainable by iterating Proposition 2(b) using Kraft’s inequality [8, Chapter 5.2].
Theorem 4 (Kraft inequality for binary trees).
Let $\ell_{1},\ldots,\ell_{k}$ be positive integers. Then, there exists a binary tree with leaves at depths $\ell_{1},\ldots,\ell_{k}$ if and only if
$\sum_{i=1}^{k}2^{-\ell_{i}}\leq 1.$
Proposition 3.
Let $P=\mathrm{conv}(P_{1},\ldots,P_{k})$ with $d(P_{i})\leq m_{i}$ for all $i$. Then
$d(P)\leq\Big\lceil\log_{2}\Big(\sum_{i=1}^{k}2^{m_{i}}\Big)\Big\rceil.$
Moreover, this is the smallest bound that can be obtained by iterating Proposition 2(b) over the polytopes $P_{i}$.
Proof.
For $k=1$, the statement is immediate. Assume $k\geq 2$, fix a binary bracketing of $\mathrm{conv}(P_{1},\ldots,P_{k})$, and let $T$ be the corresponding full binary tree.
We claim that the upper bound produced by iterating Proposition 2(b) according to $T$ is
$M_{T}=\max_{1\leq i\leq k}(m_{i}+\ell_{i}).$
We prove this by induction on the number of leaves of $T$. If $T$ has two leaves, the result is a direct consequence of Proposition 2(b). If $T$ has more than two leaves, it means that the root has two subtrees $T_{1}$ and $T_{2}$, and applying Proposition 2(b) at the root gives
$M_{T}=\max\{M_{T_{1}},M_{T_{2}}\}+1.$
By induction, each $M_{T_{j}}$ is the maximum of $m_{i}+\ell_{i}-1$ over the leaves of $T_{j}$, because their depths inside the subtree are one less than in $T$. Hence
$M_{T}=\max_{1\leq i\leq k}(m_{i}+\ell_{i}).$
Every binary tree satisfies Kraft’s inequality
$\sum_{i=1}^{k}2^{-\ell_{i}}\leq 1.$
Therefore,
$m_{i}+\ell_{i}\leq M_{T}\quad\forall i$
implies
$2^{m_{i}-M_{T}}\leq 2^{-\ell_{i}}\quad\forall i,$
and summing over $i$, we obtain
$2^{-M_{T}}\sum_{i=1}^{k}2^{m_{i}}\leq 1,$
so
$M_{T}\geq\Big\lceil\log_{2}\Big(\sum_{i=1}^{k}2^{m_{i}}\Big)\Big\rceil.$ (2)
To show that this lower bound is attained for some $T$, set
$M=\Big\lceil\log_{2}\Big(\sum_{i=1}^{k}2^{m_{i}}\Big)\Big\rceil\qquad\text{and}\qquad L_{i}=M-m_{i}.$
Therefore,
$\sum_{i=1}^{k}2^{-L_{i}}=2^{-M}\sum_{i=1}^{k}2^{m_{i}}\leq 1.$
By Kraft’s inequality, there exists a binary tree $T_{0}$ with leaves at depths $L_{1},\ldots,L_{k}$. If $T_{0}$ is not full, contract every internal node with exactly one child. This produces a full binary tree $T$ with the same labeled leaves, and the depth $\ell_{i}$ of leaf $i$ in $T$ satisfies $\ell_{i}\leq L_{i}$ for all $i$. The tree $T$ therefore determines a binary bracketing of $\mathrm{conv}(P_{1},\ldots,P_{k})$, and for this bracketing
$M_{T}=\max_{1\leq i\leq k}(m_{i}+\ell_{i})\leq\max_{1\leq i\leq k}(m_{i}+L_{i})=M.$
Combined with (2), we get $M_{T}=M$, which is then the smallest upper bound obtainable by iterating Proposition 2(b). ∎
If $m_{i}\leq m$ for all $i$ in Proposition 3, then
$d(P)\leq\Big\lceil\log_{2}\Big(\sum_{i=1}^{k}2^{m_{i}}\Big)\Big\rceil\leq\lceil\log_{2}k\rceil+m.$
We will use this simplified form to obtain depth estimates when the numbers of low-dimensional faces of a polytope are known.
For this, we also need the following concepts: $\rho(P)$ and $\rho_{2}(P)$ are the minimum number of edges and 2-faces, respectively, needed to cover all vertices of the polytope $P$; $\nu(P)$ is the size of a maximum matching of the graph $G(P)$, i.e. the largest set of edges in $G(P)$ with no common vertices; $P_{x}$ is the vertex figure of $P$ at a vertex $x$, i.e. the intersection of $P$ with a hyperplane that separates $x$ from the other vertices of $P$.
Proposition 4.
Let $f_{0},f_{1},f_{2}$ be the number of vertices, edges and 2-faces respectively of a polytope $P$. Then, we have the following depth bounds depending of each $f_{i}$ individually:
  1. (a)
    $\text{For }f_{0}:\quad d(P)\leq\lceil\log_{2}f_{0}\rceil$
  2. (b)
    $\text{For }f_{1}:\quad d(P)\leq\Big\lceil\log_{2}\Big(\Big\lfloor\frac{2f_{1}}{n}-\min\Big(n,\Big\lfloor\frac{1+\sqrt{1+8f_{1}}}{4}\Big\rfloor\Big)\Big\rfloor\Big)\Big\rceil+1$
  3. (c)
    $\text{For }f_{2}:\quad d(P)\leq\Big\lceil\log_{2}\Big(f_{2}-\binom{n}{2}+1\Big)\Big\rceil+2$
Proof.
Any polytope $P$ can be expressed as $P=\mathrm{conv}(P_{1},\ldots,P_{f_{j}})$, where $P_{i}$ are $j$-dimensional faces of $P$.
(a) If $P_{i}$ are the vertices, then $d(P_{i})=0$, and by Proposition 3 we obtain
$d(P)\leq\lceil\log_{2}f_{0}\rceil.$
(b) If $P_{i}$ are the edges of $P$, then $d(P_{i})=1$. We do not need all $f_{1}$ edges as $P$ can be written as
$P=\mathrm{conv}(P_{i_{1}},\ldots,P_{i_{\rho(P)}}),$
where $\rho(P)$ is the size of a minimum edge cover $P_{i_{1}},\ldots,P_{i_{\rho(P)}}$.
To estimate $\rho(P)$, we use standard relations between $\rho(P)$, $\nu(P)$, $f_{0}$, and $f_{1}$ (see [23, Chapter 3.1]):
$\rho(P)+\nu(P)=f_{0},\qquad\nu(P)\geq\min\big(n,\lfloor f_{0}/2\rfloor\big).$
Furthermore, the degree sum formula and the complete graph vertices bound yield
$nf_{0}\leq 2f_{1}\leq f_{0}(f_{0}-1).$
Combining these inequalities gives
$\rho(P)\leq\Big\lfloor\frac{2f_{1}}{n}-\min\Big(n,\Big\lfloor\tfrac{1+\sqrt{1+8f_{1}}}{4}\Big\rfloor\Big)\Big\rfloor,$
from which the stated bound on $d(P)$ in terms of $f_{1}$ follows.
(c) Consider the case of $2$-faces. In Section 4 we show that the depth complexity of polygons is at most $2$. Hence, similarly as (b), it suffices to bound $\rho_{2}(P)$, the minimum number of $2$-faces covering all vertices of $P$, and apply Proposition 3.
Let $P_{x}$ be the vertex figure of $P$ at a vertex $x$. Since $\dim P_{x}=n-1$, the degree sum formula implies that $P_{x}$ has at least $\binom{n}{2}$ edges. These correspond bijectively to the $2$-faces of $P$ incident to $x$. Thus, even after removing any $\binom{n}{2}-1$ 22-faces, every vertex of $P$ is still contained in at least one $2$-face. Hence $\rho_{2}(P)\leq f_{2}-\binom{n}{2}+1$, which yields
$d(P)\leq\Big\lceil\log_{2}\big(f_{2}-\binom{n}{2}+1\big)\Big\rceil+2.\qed$
3 Depth Lower Bounds
In this section, we study depth lower bounds for polytopes, obtained from their graphs and from the depth complexity of their faces.
Proposition 5.
Let $P$ be a polytope. Then for any nonempty face $F$ of $P$, we have $d(P)\geq d(F)$.
Proof.
If $d(P)=0$, there is nothing to prove. If $d(P)=1$, then $P$ is a zonotope, and any face $F$ is also a zonotope, hence $d(F)\leq 1$.
Now suppose, for induction, that the statement holds for all polytopes of depth at most m1m-1, and assume $d(P)=m$. By definition, we can write
$P=\sum_{i=1}^{q}\mathrm{conv}(P_{i1},P_{i2}),\quad d(P_{ij})<m\quad\forall i,j.$
Let $F$ be a face of $P$. Then $F=\sum_{i=1}^{q}F_{i}$, where each $F_{i}$ is a face of $\mathrm{conv}(P_{i1},P_{i2})$. This implies that $F_{i}$ can be written as $F_{i}=\mathrm{conv}(F_{i1},F_{i2})$, where $F_{ij}$ is a (possibly empty) face of $P_{ij}$.
By the induction hypothesis, whenever $F_{ij}\neq\emptyset$, we have $d(F_{ij})<m$. If both $F_{i1}$ and $F_{i2}$ are nonempty, then by Proposition 2(b) we have $d(F_{i})\leq m$. If exactly one of $F_{i1},F_{i2}$ is nonempty, then $F_{i}$ is equal to that nonempty face, and hence $d(F_{i})\leq m$.
Therefore, in all cases each $F_{i}$ satisfies $d(F_{i})\leq m$. Finally, since $F=\sum_{i=1}^{q}F_{i}$, Proposition 2(a) implies $d(F)\leq m=d(P)$. ∎
To obtain a lower bound from complete subgraphs contained in $G(P)$, we first state the following lemma.
Lemma 1.
If the graph of a polytope $G(P)$ contains a complete subgraph with $k\geq 3$ vertices, and $P$ can be decomposed as $P=\sum_{i=1}^{q}P_{i}$, then at least one of $G(P_{i})$ also contains a complete subgraph with $k$ vertices.
Proof.
Any edge $e=[x_{1},x_{2}]$ of $P$ can be written as $e=\sum_{i=1}^{q}e_{i}$, where each $e_{i}$ is either a vertex or an edge of $P_{i}$. If $e_{i}$ is a vertex, we regard it as a degenerate edge of length zero, so we can say that each edge of $P$ decomposes as a sum of edges of the $P_{i}$.
Now let $u,v,w$ be three vertices in the complete subgraph of $G(P)$ (so $[u,v],[u,w],[v,w]$ are edges of $P$). Write
$u,v]=\sum_{i=1}^{q}[u_{i},v_{i}],\quad[u,w]=\sum_{i=1}^{q}[u_{i},w_{i}],\quad[v,w]=\sum_{i=1}^{q}[v_{i},w_{i}],$
where $u=\sum_{i}u_{i}$, $v=\sum_{i}v_{i}$, $w=\sum_{i}w_{i}$ with $u_{i},v_{i},w_{i}$ vertices of $P_{i}$.
The edges $u_{i},v_{i}],[u_{i},w_{i}],[v_{i},w_{i}]$ are parallel to $[u,v],[u,w],[v,w]$ respectively, since otherwise their Minkowski sum would yield a polytope of dimension at least two, contradicting the fact that the corresponding sum is an edge. This implies that each triangle $u_{i},v_{i},w_{i}]$ is homothetic to the triangle $[u,v,w]$, and therefore all edge length ratios are preserved
$\frac{|u_{i}-v_{i}|}{|u-v|}=\frac{|u_{i}-w_{i}|}{|u-w|}=\frac{|v_{i}-w_{i}|}{|v-w|}.$
Note that some triangles $u_{i},v_{i},w_{i}]$ may be degenerate (if two or all three vertices coincide). In this case, the triangle is still homothetic to $[u,v,w]$ with homothety factor 0, so the ratio identities above remain valid.
As the triangle $[u,v,w]$ is not degenerate, at least one index $j$ has these ratios nonzero, so $u_{j},v_{j},w_{j}$ are distinct vertices forming a triangle in $G(P_{j})$.
For any other vertex $z$ in the clique of $G(P)$, applying the same argument to the triangles $[u,v,z]$ and $[u,w,z]$ shows that the same index $j$ yields non-degenerate homothetic triangles $u_{j},v_{j},z_{j}]$ and $u_{j},w_{j},z_{j}]$, with $z_{j}$ distinct from $u_{j},v_{j},w_{j}$ and adjacent to them. Repeating this for every vertex in the clique, we conclude that $G(P_{j})$ contains a complete subgraph with $k$ vertices. ∎
Theorem 5.
If the graph of a polytope $G(P)$ contains a complete subgraph with $k\geq 3$ vertices, then $d(P)\geq\lceil\log_{2}k\rceil$.
Proof.
Suppose a subgraph of $G(P)$ is complete and contains $k=3$ or $k=4$ vertices. If we assume $d(P)=1$, then $P=\sum_{i=1}^{k}P_{i}$, where each $P_{i}$ is a segment. This contradicts Lemma 1, which implies that at least one $P_{i}$ must include $k$ vertices. Therefore, we conclude $d(P)\geq 2$.
For the sake of induction, let’s assume that the result holds for all cases up to k1k-1. Now, consider that $G(P)$ includes a complete subgraph consisting of $k$ vertices. By definition, we can express $P$ as
$P=\sum_{i=1}^{q}\mathrm{conv}(P_{i},Q_{i}),\quad\textrm{where }d(P_{i}),d(Q_{i})<d(P).$
According to Lemma 1, there exists an index $i$ for which $G(\mathrm{conv}(P_{i},Q_{i}))$ also contains a complete subgraph $K$ with $k$ vertices. Without loss of generality, we can assume that $P_{i}$ contains at least $\lceil\frac{k}{2}\rceil$ vertices of $K$.
If two vertices $x,y$ of $\mathrm{conv}(P_{i},Q_{i})$ span an edge, then they also span an edge in $P_{i}$ if $x,y\in P_{i}$. Hence $G(P_{i})$ contains not only at least $\lceil\frac{k}{2}\rceil$ vertices of $K$ but also the complete subgraph induced by them. By the induction hypothesis we obtain
$d(P)-1\geq d(P_{i})\geq\Big\lceil\log_{2}\Big\lceil\frac{k}{2}\Big\rceil\Big\rceil=\big\lceil\log_{2}k\big\rceil-1,$
from which it follows $d(P)\geq\lceil\log_{2}k\rceil$. ∎
4 Depth of Polytopes
We begin by computing the depth complexity of polygons, whose depth is at most 2.
Polygons. $d(P)=1$ if $P$ is a zonotope, otherwise $d(P)=2$.
Proof.
If $P$ is a zonotope, then $d(P)=1$. If $P$ is a triangle, then $d(P)=2$ by Proposition 4(a), since a triangle is not a zonotope.
Now suppose $P$ is neither a zonotope nor a triangle. Then, $P$ can be decomposed as $P=\sum_{i=1}^{k}P_{i},$ where each $P_{i}$ is either a zonotope or a triangle [15, Chapter 15.1]. By Proposition 2(a), we obtain $d(P)=2$. ∎
A pyramid is a polytope of the form $P=\mathrm{conv}(Q,x)$, where $Q$ is the basis and $x\in\mathbb{R}^{n}$ such that $x\not\in\text{aff }Q$, the affine hull of $Q$.
Pyramids. $d(Q)\leq d(P)\leq d(Q)+1$, where $Q$ is a basis of $P$.
Proof.
From Proposition 2(b) we obtain $d(P)\leq d(Q)+1$. Since $Q$ is a face of $P$, Proposition 5 gives $d(Q)\leq d(P)$. ∎
A bipyramid is a polytope of the form $P=\mathrm{conv}(Q,x_{1},x_{2})$, where $Q$ (with $\dim Q\geq 1$) is the base, and $x_{1},x_{2}\notin\text{aff }Q$ are the apices, such that the segment $x_{1},x_{2}]$ intersects the relative interior of $Q$.
Bipyramids. $d(Q)\leq d(P)\leq d(Q)+1$, where $Q$ is the basis of $P$.
Proof.
The upper bound $d(P)\leq d(Q)+1$ follows from Proposition 2(b). For the lower bound, let $H$ be a hyperplane containing aff $Q$ but not $x_{1},x_{2}$. Projecting $P$ onto $H$ along the line through the apices gives $Q$, hence by Proposition 2(c) we have $d(Q)\leq d(P)$. ∎
Cartesian product. $d(P_{1}\times P_{2})=\max\{d(P_{1}),d(P_{2})\}.$
Proof.
Let
$P_{1}^{\prime}=P_{1}\times\{0\},\qquad P_{2}^{\prime}=\{0\}\times P_{2}.$
Then $P_{1}\times P_{2}=P_{1}^{\prime}+P_{2}^{\prime}$. Since $P_{1}^{\prime}$ and $P_{2}^{\prime}$ are affinely equivalent to $P_{1}$ and $P_{2}$, respectively, Proposition 2(a, c) gives
$d(P_{1}\times P_{2})\leq\max\{d(P_{1}),d(P_{2})\}.$
For the reverse inequality, choose vertices $x_{i}$ of $P_{i}$. Then $P_{1}\times\{x_{2}\}$ and $\{x_{1}\}\times P_{2}$ are faces of $P_{1}\times P_{2}$ affinely equivalent to $P_{1}$ and $P_{2}$. By Proposition 5,
$d(P_{i})\leq d(P_{1}\times P_{2}),\ \ i=1,2.$
Hence
$\max\{d(P_{1}),d(P_{2})\}\leq d(P_{1}\times P_{2}),$
which proves the claim. ∎
A prism is a polytope of the form $P=\mathrm{conv}(Q,x+Q)$, where $Q$ is called the base of $P$ and the translation $x+Q\notin\text{aff }Q$.
Prisms. $d(P)=d(Q)$, where $Q$ is the base of $P$.
Proof.
A prism with base $Q$ can be written as $P=Q\times[0,1]$ up to affine equivalence. The claim follows from the cartesian product depth, since $d([0,1])=1$. ∎
The cross-polytope is polytope of the form $P=\mathrm{conv}(\pm e_{1},\ldots,\pm e_{n})$, where $e_{i}$ are the standard basis vectors of $\mathbb{R}^{n}$.
Cross-polytopes. $d(P)=\lceil\log_{2}n\rceil$ for $n\geq 2$.
Proof.
The vertices $e_{1},\ldots,e_{n}$ form a clique in the graph of $P$, so by Theorem 5,
$d(P)\geq\lceil\log_{2}n\rceil.$
For the upper bound, proceed by induction on $n$. The case $n=2$ is a square, hence a zonotope, so $d(P)=1$.
Assume $n\geq 3$ and write $n=a+b$ with $a=\lfloor n/2\rfloor$ and $b=\lceil n/2\rceil$. Identifying $\mathbb{R}^{n}=\mathbb{R}^{a}\times\mathbb{R}^{b}$, we have
$P=\mathrm{conv}(P_{a}\times\{0\},\;\{0\}\times P_{b}),$
where $P_{a},P_{b}$ are the cross-polytopes of dimension $a$ and $b$, respectively. Therefore, by Proposition 2(b, c), the cartesian product depth, and the induction hypothesis,
$d(P)$ $\leq\max\{d(P_{a}\times\{0\}),d(\{0\}\times P_{b})\}+1$
$=\max\{d(P_{a}),d(P_{b})\}+1$
$=\lceil\log_{2}b\rceil+1$
$\leq\lceil\log_{2}n\rceil.$
Hence $d(P)=\lceil\log_{2}n\rceil$. ∎
A 2-neighborly polytope $P$ is a polytope where the segment between any two vertices is an edge of $P$.
2-neighborly polytopes. $d(P)=\lceil\log_{2}k\rceil$, where $k$ is the number of vertices of $P$.
Proof.
Since $G(P)$ is complete, this follows from Proposition 4(a) and Theorem 5. ∎
As simplices are 2-neighborly we obtain directly their depth complexity.
Simplices. $d(P)=\lceil\log_{2}(n+1)\rceil$.
Using Theorem 3, we obtain a geometric proof of Theorem 1 for ReLU networks: the Newton polytope of $\max\{x_{1},\ldots,x_{n},0\}$ is an $n$-simplex, so the computed depth for simplices recovers the same depth bound as Theorem 1.
The join of polytopes is a polytope of the form $P_{1}*\cdots*P_{k}=\mathrm{conv}(P_{1},\ldots,P_{k})$, where the $P_{i}$ lie in pairwise skew affine subspaces.
Join of $(2^{m_{i}}-1)$-simplices. $d(P)=\Big\lceil\log_{2}\Big(\sum_{i=1}^{k}2^{m_{i}}\Big)\Big\rceil$, for $P=P_{1}*\cdots*P_{k}$, where each $P_{i}$ is a $(2^{m_{i}}-1)$-simplex.
Proof.
For each $i$, the simplex $P_{i}$ has depth $d(P_{i})=m_{i}$. Since $P=\mathrm{conv}(P_{1},\ldots,P_{k})$, Proposition 3 gives
$d(P)\leq\Big\lceil\log_{2}\Big(\sum_{i=1}^{k}2^{m_{i}}\Big)\Big\rceil.$
On the other hand, the graph of $P$ is complete on $\sum_{i=1}^{k}2^{m_{i}}$ vertices, since every pair of vertices from different $P_{i}$ is adjacent. Therefore, Theorem 5 yields
$d(P)\geq\Big\lceil\log_{2}\Big(\sum_{i=1}^{k}2^{m_{i}}\Big)\Big\rceil,$
proving the claim. ∎
The depth computation for join of simplices shows that for every vector $(m_{1},\ldots,m_{k})\in\mathbb{N}^{k}$, there are polytopes attaining exactly the upper bound of Proposition 3.
A cyclic polytope $C_{n}(k),k>n,$ is a polytope of the form
$C_{n}(k)=\mathrm{conv}(\gamma(t_{1}),\ldots,\gamma(t_{k})),$
where $t_{1}<\ldots<t_{k}$ are real numbers and $\gamma(t)=(t,t^{2},\ldots,t^{n})$. For $n\geq 4$, cyclic polytopes are 2-neighborly, therefore we obtain directly their depth complexity.
Cyclic polytopes. $d(C_{n}(k))=\lceil\log_{2}k\rceil$ for $n\geq 4$.
Since $d(P)\leq d_{0}(P)$, where $d_{0}(P)$ is the minimum ICNN depth required to represent $\mathcal{F}P$, the depth of cyclic polytopes for $n\geq 4$ shows that ICNNs do not satisfy a result like Theorem 1. Unlike general ReLU networks, no fixed depth suffices to represent $\mathcal{F}C_{n}(k)$ as $k$ grows.
For $n=2$, polygons were shown to have bounded depth $d(P)\leq 2$, while the case $n=3$ remains open. We show that bipyramids with a triangular base provide examples of polyhedra of depth 3. This demonstrates that depth complexity in dimension $n=3$ behaves differently from the case $n=2$.
Before computing the depth complexity of triangular bipyramids, we first introduce a simple tool for indecomposable polytopes.
Two polytopes, $P$ and $Q$, are said to be positively homothetic, if $P=\lambda Q+w$ for some $\lambda>0$ and $w\in\mathbb{R}^{n}$. A polytope $P$ is said to be indecomposable if any decomposition $P=\sum_{i=1}^{k}P_{i}$ is only possible when $P_{i}$ is positively homothetic to $P$ for all $i=1,\ldots,k$.
Lemma 2.
If $P$ is an indecomposable polytope, then there exist polytopes $P_{1},P_{2}$ such that $P=\mathrm{conv}(P_{1},P_{2})$ and $d(P)=\max\{d(P_{1}),d(P_{2})\}+1$.
Proof.
By definition, if $d(P)=m$, then $P$ can be written as
$P=\sum_{i=1}^{q}\mathrm{conv}(P_{i1},P_{i2}),\quad d(P_{ij})<m\quad\forall i,j,$
where $m$ is the smallest integer for which such a representation exists. Hence there must exist an index $i$ such that $\max\{d(P_{i1}),d(P_{i2})\}+1=d(P)$.
Since $P$ is indecomposable, there are $\lambda_{i}>0$ and $w_{i}\in\mathbb{R}^{n}$ where $P$ satisfies
$P=\lambda_{i}\mathrm{conv}(P_{i1},P_{i2})+w_{i}=\mathrm{conv}(\lambda_{i}P_{i1}+w_{i},\lambda_{i}P_{i2}+w_{i}).$
Each $\lambda_{i}P_{ij}+w_{i}$ is an affine image of $P_{ij}$ under an invertible map. By Proposition 2(c), affine equivalence preserves depth, so
$d(P)=\max\{d(P_{i1}),d(P_{i2})\}+1=\max\{d(\lambda_{i}P_{i1}+w_{i}),d(\lambda_{i}P_{i2}+w_{i})\}+1.\qed$
Triangular bipyramids in $\mathbb{R}^{3}$. $d(P)=3$.
Proof.
A bipyramid with a 2-depth basis satisfies $2\leq d(P)\leq 3$. Assume, for contradiction, that $d(P)=2$. As $P$ is indecomposable [15, Chapter 15.1], Lemma 2 implies that there exist $P_{1},P_{2}$ such that $P=\mathrm{conv}(P_{1},P_{2})$ and $\max\{d(P_{1}),d(P_{2})\}+1=2$. Without loss of generality, suppose $P_{1}$ is a zonotope containing at least three vertices of $P$.
There are three cases: $P_{1}$ contains two base vertices and an apex; or the three base vertices; or one base vertex and the two apices. In each case $P_{1}$ is not centrally symmetric, contradicting that it is a zonotope. Hence $d(P)\neq 2$ and so $d(P)=3$. ∎
We conclude with a complementary construction: whereas for cyclic polytopes the depth grows with the number of vertices, here we obtain polytopes with arbitrarily many vertices at fixed depth.
Theorem 6.
For any $n\geq 5$ and $m\in\mathbb{N}$, there exists a family of $n$-polytopes with arbitrarily many vertices and depth $m$.
Proof.
Let $g_{i}=[\mathbf{0},b_{i}],i=1,\ldots,p$, represent line segments from the origin with $b_{1},\ldots,b_{p}$ denoting $p\geq n$ points in general position. The zonotope $Z_{p}=\sum_{i=1}^{p}g_{i}$ has depth 1 and $v_{p}=2\sum_{i=0}^{n-1}\binom{p-1}{i}$ vertices [26], which grows with $p$.
For any $m\geq 2$, let $H$ be a supporting hyperplane of a vertex $x$ of $Z_{p}$, and take a polytope $P\subset H$ with $\dim P<n$ and $d(P)=m$ (e.g. a simplex or cyclic polytope).
Then $P+Z_{p}$ satisfies $d(P+Z_{p})\leq m$ by Proposition 2(a). On the other hand, the face of $P+Z_{p}$ in the direction of $x$ is a translate of $P$, hence by Proposition 5, $m=d(P)\leq d(P+Z_{p})$. Thus $P+Z_{p}$ forms a $m$-depth family whose number of vertices increases with $p$. ∎
Similar constructions exist for $n\leq 4$ and $m\leq 3$, but a full generalization of Theorem 6 for $n=3,4$ and arbitrary $m$ would require a better understanding of depth complexity in dimension $3$.